An engineer is designing a process for a new transistor. She uses a vacuum chamber to bombard a thin layer of silicon with ions of phosphorus, each of mass mP = 5.18 × 10-26 kg. The phosphorus ions are doubly ionized, with each phosphorus ion lacking two electrons. The ions start at rest at one end of the vacuum chamber and are accelerated by an electric field over a distance of re = 35 cm before they strike the silicon layer with velocity vP = 195 m/s.

A. Enter an expression for the potential difference ΔV, in volts, between the initial and final points across the vacuum chamber.

B. Calculate the average electric field strength E, in volts per meters, across the vacuum chamber.

C. Calculate the average electric force F, in newtons, that the electric field exerts on each phosphorus ions.

&

Suppose the electric field between two conducting plates has a strength of 41 × 103 V/m.

How far apart, in meters, are the two conducting plates if their potential difference is 0.145 kV?

&

The voltage across a membrane forming a cell wall is 75 mV and the membrane is 7.75 nm thick.

What is the electric field strength in the cell wall, in volts per meter?
Answer #1

mo solution. Given that velocity up = 195 m/s Mass mp= 5-18 x 10-26 kg re= 35cm @ Potential difference Av= -1 mpvp Average ei hope this answer will be helpful to you

please like this answer

thank you